there are two planes fly in two directions one is flying for east direction with the rate of 300kph and other is for west direction with the rate of 360kph. in how many minutes if they are 660km. apart.
please show the solution.thank you.

Answer by stanbon(75887)

there are two planes fly in two directions one is flying for east direction with the rate of 300kph and other is for west direction with the rate of 360kph. in how many minutes if they are 660km. apart.
------------------
They are separating at the rate of 300 + 360 = 660 km/hr
----
time = distance/rate
time = 660/660 = 1 hr = 60 minutes
======================================
Cheers,
Stan H.
======================